# Learning Multi-frame and Monocular Prior for Estimating Geometry in Dynamic Scenes

### v1

@@ 1 Introduction
However, the inherent drawback of the concurrent models is that they process only a pair of frames at once, and extending the number of frames is non-trivial in their Siamese architecture. This poses significant limitations for processing complex dynamic scenes that require observing multiple frames beyond the pairs, and the models demonstrate sub-optimal performance, as depicted in Figure 1. While existing methods mitigate the errors by accumulating pairwise estimates for multiple frames through the global optimization, they inevitably are computationally heavy and prone to errors, akin to the classical optimization-based approaches.
@@
We provide the details of our method in Section 3, the preliminaries of the Siamese architecture and our specific designs to address the problem. Then, we perform experiments benchmarking the quality of 4D geometry estimation in comparison with state-of-the-art baselines in Section 4, where MMP achieves significant improvement in the feed-forward prediction quality.
@@ 3 Method
In this section, we provide the details of our architecture design for predicting pointmaps given a set of video frames. To begin, we review the baseline Siamese architecture in Section 3.1, based on which we design a new architecture for our method. Then, we introduce the trajectory encoder in Section 3.2, the key component of our method, which enables processing multiple frames beyond the limitation of the baseline. Finally, in Section 3.3, we describe the feed-forward refinement technique in our method.
As for the data notation, we denote scalars using normal letters, and tensors using bold letters with a superscript denoting frame indices. For example, an input RGB video frame is $\mathbf{I}^{i}\in\mathbb{R}^{U\times V\times 3}$, where $U\times V$ is the resolution, and a frame tokenization is $\mathbf{F}^{i}\in\mathbb{R}^{N\times D}$, where $N=\frac{U}{P}\times\frac{V}{P}$ with the patch size $P$ and the embedding dimension $D$. Tensors can be indexed, such as $\mathbf{F}^{i}(n)\in\mathbb{R}^{D}$, where $\mathbf{F}^{i}\equiv[\mathbf{F}^{i}(1),...,\mathbf{F}^{i}(N)]$. Finally, when emphasizing that a feature or data for frame $i$ is conditioned on the frame $j$, we use the superscript $i|j$, such as the pointmap output $\mathbf{X}^{i|j}\in\mathbb{R}^{U\times V\times 3}$, which we frequently use in Section 3.1.
@@ 3.1 Pair-wise Siamese architecture
Given a pair of frames $(\mathbf{I}^{i},\mathbf{I}^{j})$, the Siamese architecture aims to predict a pointmap: the ego pointcloud $\mathbf{X}^{i|j}$ which represents the 3D coordinate of $\mathbf{I}^{i}$, and the target pointcloud $\mathbf{Y}^{j|i}$ which represents the 3D coordinate of $\mathbf{I}^{j}$ following the camera view of $\mathbf{I}^{i}$, predicted by two separate decoders. Specifically, concurrent models (Wang et al., 2024; Zhang et al., 2025) employ transformer blocks with relative position embedding as the decoder, which process the ego tokens $\mathbf{E}_{l}^{i|j}\in\mathbb{R}^{N\times D}$, and the target tokens $\mathbf{T}_{l}^{i|j}\in\mathbb{R}^{N\times D}$, where $l\in\{1,...,L\}$ is the transformer block index. The initial tokens ($l=0$) are ${\mathbf{F}^{i}:=\mathtt{Tokenization}(\mathbf{I}^{i})}$, i.e.$\mathbf{E}_{0}^{i|j}:=\mathbf{F}^{i}$ and $\mathbf{T}_{0}^{j|i}:=\mathbf{F}^{j}$.
In each transformer block (Figure 2b), the cross-attention $\mathtt{CA}(\cdot;\cdot)$, placed next to the self-attention $\mathtt{SA}(\cdot)$, conveys information between the ego and the target tokens, followed by the $\mathtt{MLP}(\cdot)$ layer producing the output of a block,
@@
For most use cases, pair-wise models are executed twice, under the original and a swapped order of the input frames, producing $\{\mathbf{X}^{i|j},\mathbf{X}^{j|i},\mathbf{Y}^{j|i},\mathbf{Y}^{i|j}\}$, which enables downstream tasks, such as 2-view geometry, estimating camera intrinsics and pose, etc. When processing a greater number of frames $W>2$, inference is performed over all combinations, e.g., for all $i\neq j$, $i\in\{1,...,W\}$ and $j\in\{1,...,W\}$. However, the pair-wise architecture is limited to process complex dynamic scenes, and the feed-forward performance is often sub-optimal, as we find in Section 4.3.
@@ 3.2 Trajectory encoder
In this section, we describe our method to jointly process multiple frames (i.e.$W>2$) to predict dynamic pointmaps. To be specific, we enable it using the proposed trajectory encoder module, which collects the tokens in the same spatial index over the frames, then encode the inter-frame dynamics back to each token.
Without loss of generality, let us consider the frame $\mathbf{I}^{W}$, paired with others $\{\mathbf{I}^{1},...,\mathbf{I}^{W-1}\}$ and their corresponding tokens within the intermediate cross-attention stage of the decoder blocks in Equations 1 and 3,
@@
Intuitively, gathering from a same spatial index, e.g., a stack of tokens $[\tilde{\mathbf{T}}_{l}^{W|1}(n),...,\tilde{\mathbf{T}}_{l}^{W|W-1}(n)]\in\mathbb{R}^{W\times D}$ by indexing each element in Equation 7, can represent the spatio-temporal dynamics of the patch region represented by $\mathbf{F}^{W}(n)$. Therefore, projecting this feature onto each token of Equations 7 and 8 can encode the dynamics. Specifically, we apply an attention mechanism [^1] We adjust the relative position embedding to encode a spatial index with the size $D/2$, and a time index with the size $D/2$. with causal masks to implement the function, coined trajectory attention $\mathtt{TA}(\cdot;\cdot)$,
@@ 3.3 Feed-forward refinement
Although MMP architecture does not constrain the number of frames $W$, the finite memory of the system can pose a practical limit. When processing tens or hundreds of frames as the prediction horizon, a joint processing of whole frames can be impossible. In order to overcome the limitation, we introduce a feed-forward refinement technique to deal with prediction horizon beyond a chosen $W$. Specifically, when processing the tokens of an extra frame, e.g.$\tilde{\mathbf{E}}_{l}^{i|j}$, where $i\leq W$ and $j>W$, we exploit the pre-computed key and value tensors of $\tilde{\mathbf{E}}_{l}^{i|\{k<W\}}$, which we illustrate in Figure 3. Since we train the model with the causal attention masking applied to the trajectory attention, these key and value tensors remain equivalent to the case where a larger input size were considered to include the extra frame.
@@ 4 Experiment
In this section, we present the experimental details and compare MMP to state-of-the-art baselines. In Sections 4.1 and 4.2, we provide the training details, the data processing, and the evaluation protocols. Then, we experiment, in Section 4.3, the feed-forward prediction of pointmaps, and the ablation study in Section 4.4.
@@ 4.1 Training details
Despite our design to maintain the strong 3D prior of the pre-trained model (Wang et al., 2024), the synthetic scenes in the training dataset can cause a distribution shift in visual texture. Therefore, we test the trade-off between different training schedules for mixing the synthetic and the real frames, then choose the default setting that demonstrates a balanced performance (see Section 4.4 for more details). Our default setting trains MMP for 30 epochs using the AdamW optimizer (Loshchilov and Hutter, 2019) with 20k clips of length $W=5$ per epoch, the mini-batch size 16, and the learning rate $1\times 10^{-4}$. We sample the clips from real scenes for the first 5 epochs, then employ synthetic scenes for the rest of the training steps.
@@ 4.2 Evaluation details
To evaluate the feed-forward predictions (Section 4.3), we employ 3 different test datasets covering dynamic scenes: Point Odyssey (Zheng et al., 2023), Sintel (Butler et al., 2012), and iPhone dataset (Gao et al., 2022). Point Odyssey and Sintel are synthesized scenes generated using 3D rendering engines (Zheng et al., 2023; Butler et al., 2012), and iPhone dataset covers real scenes captured using a synchronized set of camera, lidar, and IMU sensors (Gao et al., 2022). For each scene, we consider overlapping slices of 12 frames as the evaluation samples. [^2] We also downsample iPhone dataset (Gao et al., 2022) to 3fps to promote a larger motion. We measure the regression accuracy of the pointmaps predicted by MMP and the baselines: DUSt3R (Zhang et al., 2025), Robust-CVD (Kopf et al., 2021), CasualSAM (Zhang et al., 2022), and state-of-the-art MonST3R (Zhang et al., 2025). Specifically, using a strided sampling, we experiment with $W=2$ (stride 6), $W=4$ (stride 3), and $W=6$ (stride 2) for inference. As for the metric, we employ the scale and shift invariant error provided by the open source repository of MonST3R (Zhang et al., 2025) and report the median error in the target pointclouds per setting: M@2, M@4, and M@6 in Table 1.
@@ 4.3 Feed-forward pointmap prediction
In this section, we experiment the feed-forward pointmap prediction by MMP. In Table 1, we quantitatively compare the quality of pointmap regression by MMP and the baselines: DUSt3R (Zhang et al., 2025), Robust-CVD (Kopf et al., 2021), CasualSAM (Zhang et al., 2022), and MonST3R (Zhang et al., 2025). Next, we provide the visualization of the pointmaps produced by MMP in Figure 4, executed on DAVIS video frames (Perazzi et al., 2016).
To begin with, we find MMP can outperform the strongest feed-forward baseline, MonST3R (Zhang et al., 2025), e.g., 15.1% improvement M@6 1.772 (MonST3R (Zhang et al., 2025)) $\to$ 1.504 (MMP) on iPhone dataset (Gao et al., 2022) in Table 1. While our method is trained on the same data distribution as the baseline, an enhanced performance is observed even under a pair-wise inference (i.e., M@2). This supports the significance of the trajectory encoder employed in our method, which facilitates learning useful representation for predicting accurate pointmaps. MMP can consistently improve the quality of dynamic pointmaps compared to the baselines in various scenarios covering synthetic and real video scenes. We also note that our method can demonstrate the results that are more robust over various strides, (e.g., comparing to MonST3R (Zhang et al., 2025) in Sintel (Butler et al., 2012): 1.374 $\to$ 1.298 M@2, 1.411 $\to$ 1.288 M@4, and 1.433 $\to$ 1.287 M@6), which we attribute to the dynamics modeling enabled by our method.
@@
From the qualitative study in Figure 4, we find our method tends to demonstrate more accurate pointmaps over the frames, e.g., the background objects and the scene are consistently depicted, comparing the regions indicated by red boxes (MonST3R (Zhang et al., 2025)) and the blue boxes (MMP), which reveals the efficacy of our method in complex dynamic scenes.
@@ 4.4 Ablation study
We further study the effect of applying different schedules for training synthetic and real scenes by MMP in Table 3, which compares 4 different training strategies: synthetic only (i.e., synthetic scenes for 30 epochs), joint training (i.e., mixed data for 30 epochs), synthetic then real (i.e., synthetic scenes for the first 25 epochs, then real scenes for the rest 5 epochs), and real then synthetic (the default setting). While there exist trade-offs in the performances over the datasets, we choose the real then synthetic schedule as our final design, which can demonstrate a balanced performance.
@@ 5.1 Extreme case
However, in practice, including the scenarios represented by our evaluation, there exists an extreme case where $\mathbf{I}^{i}$ and $\mathbf{I}^{k}$ are completely non-overlapping, so that the pair-wise model assigns an erroneous estimate of $\mathtt{Pr}(\mathbf{Y}^{i|k})$, which can induce significant failure modes of estimating the joint density. Even if the global optimization is employed, depending on the sampling strategy, there is a potential extreme case that the connectivity becomes independent. To prevent the case, a sophisticated hyperparameter engineering would be required. Since MMP can relax this constraint up to $W$ frames and beyond (with the feed-forward refinement), it can learn the pointmap distribution that is more close to the true nature of the dynamic scenes. For example, the intriguing tendency of MMP in Table 1, being robust to the evaluation stride can be attributed to a more accurate estimation of the joint density over a set of frames.
@@ 5.2 Limitation
Despite the promising results demonstrated by MMP, the scarcity of 4D dyanamic scenes can hinder the generalization performance. To mitigate the distribution shifts, we employ the scheduled training to maintain the visual texture prior in the pre-trained model. However, since we still observe trade-offs in the performance, as shown in Table 3, designing new training datasets, self-supervised learning with unlabeled data, or an objective functions robust to the distribution shift for 4D geometry estimation can be interesting future directions. It is also worth noting that we focus on the realistic scenarios where the observation is captured by a monocular video camera, rather than multiple synchronized cameras capturing one scene. Although it would be straightforward to apply MMP for the synchronized cameras, we believe that there is a room to exploit useful properties, such as epipolar geometry (Hartley and Zisserman, 2003) of the synchronized cameras, which is another interesting future direction.

### v2

@@ 1 Introduction
However, the inherent drawback of the concurrent models is that they process only a pair of frames at once, and extending the number of frames is non-trivial in their Siamese architecture. This poses significant limitations for processing complex dynamic scenes that require observing multiple frames beyond the pairs, and the models demonstrate sub-optimal performance, as depicted in Figure˜1. While existing methods mitigate the errors by accumulating pairwise estimates for multiple frames through the global optimization, they inevitably are computationally heavy and prone to errors, akin to the classical optimization-based approaches.
@@
We provide the details of our method in Section˜3, the preliminaries of the Siamese architecture and our specific designs to address the problem. Then, we perform experiments benchmarking the quality of 4D geometry estimation in comparison with state-of-the-art baselines in Section˜4, where MMP achieves significant improvement in the feed-forward prediction quality.
@@ 3 Method
In this section, we provide the details of our architecture design for predicting pointmaps given a set of video frames. To begin, we review the baseline Siamese architecture in Section˜3.1, based on which we design a new architecture for our method. Then, we introduce the trajectory encoder in Section˜3.2, the key component of our method, which enables processing multiple frames beyond the limitation of the baseline. Finally, in Section˜3.3, we describe the feed-forward refinement technique in our method.
As for the data notation, we denote scalars using normal letters, and tensors using bold letters with a superscript denoting frame indices. For example, an input RGB video frame is $\mathbf{I}^{i}\in\mathbb{R}^{U\times V\times 3}$, where $U\times V$ is the resolution, and a frame tokenization is $\mathbf{F}^{i}\in\mathbb{R}^{N\times D}$, where $N=\frac{U}{P}\times\frac{V}{P}$ with the patch size $P$ and the embedding dimension $D$. Tensors can be indexed, such as $\mathbf{F}^{i}(n)\in\mathbb{R}^{D}$, where $\mathbf{F}^{i}\equiv[\mathbf{F}^{i}(1),...,\mathbf{F}^{i}(N)]$. Finally, when emphasizing that a feature or data for frame $i$ is conditioned on the frame $j$, we use the superscript $i|j$, such as the pointmap output $\mathbf{X}^{i|j}\in\mathbb{R}^{U\times V\times 3}$, which we frequently use in Section˜3.1.
@@ 3.1 Pair-wise Siamese architecture
Given a pair of frames $(\mathbf{I}^{i},\mathbf{I}^{j})$, the Siamese architecture aims to predict a pointmap: the ego pointcloud $\mathbf{X}^{i|j}$ which represents the 3D coordinate of $\mathbf{I}^{i}$, and the target pointcloud $\mathbf{Y}^{j|i}$ which represents the 3D coordinate of $\mathbf{I}^{j}$ following the camera view of $\mathbf{I}^{i}$, predicted by two separate decoders. Specifically, concurrent models (Wang et al., 2024; Zhang et al., 2025) employ transformer blocks with relative position embedding as the decoder, which process the ego tokens $\mathbf{E}_{l}^{i|j}\in\mathbb{R}^{N\times D}$, and the target tokens $\mathbf{T}_{l}^{i|j}\in\mathbb{R}^{N\times D}$, where $l\in\{1,...,L\}$ is the transformer block index. The initial tokens ($l=0$) are ${\mathbf{F}^{i}:=\mathtt{Tokenization}(\mathbf{I}^{i})}$, i.e., $\mathbf{E}_{0}^{i|j}:=\mathbf{F}^{i}$ and $\mathbf{T}_{0}^{j|i}:=\mathbf{F}^{j}$.
In each transformer block (Figure˜2b), the cross-attention $\mathtt{CA}(\cdot;\cdot)$, placed next to the self-attention $\mathtt{SA}(\cdot)$, conveys information between the ego and the target tokens, followed by the $\mathtt{MLP}(\cdot)$ layer producing the output of a block,
@@
For most use cases, pair-wise models are executed twice, under the original and a swapped order of the input frames, producing $\{\mathbf{X}^{i|j},\mathbf{X}^{j|i},\mathbf{Y}^{j|i},\mathbf{Y}^{i|j}\}$, which enables downstream tasks, such as 2-view geometry, estimating camera intrinsics and pose, etc. When processing a greater number of frames $W>2$, inference is performed over all combinations, e.g., for all $i\neq j$, $i\in\{1,...,W\}$ and $j\in\{1,...,W\}$. However, the pair-wise architecture is limited to process complex dynamic scenes, and the feed-forward performance is often sub-optimal, as we find in Section˜4.3.
@@ 3.2 Trajectory encoder
In this section, we describe our method to jointly process multiple frames (i.e., $W>2$) to predict dynamic pointmaps. To be specific, we enable it using the proposed trajectory encoder module, which collects the tokens in the same spatial index over the frames, then encode the inter-frame dynamics back to each token.
Without loss of generality, let us consider the frame $\mathbf{I}^{W}$, paired with others $\{\mathbf{I}^{1},...,\mathbf{I}^{W-1}\}$ and their corresponding tokens within the intermediate cross-attention stage of the decoder blocks in Equations˜1 and 3,
@@
Intuitively, gathering from a same spatial index, e.g., a stack of tokens $[\tilde{\mathbf{T}}_{l}^{W|1}(n),...,\tilde{\mathbf{T}}_{l}^{W|W-1}(n)]\in\mathbb{R}^{W\times D}$ by indexing each element in Equation˜7, can represent the spatio-temporal dynamics of the patch region represented by $\mathbf{F}^{W}(n)$. Therefore, projecting this feature onto each token of Equations˜7 and 8 can encode the dynamics. Specifically, we apply an attention mechanism [^1] We adjust the relative position embedding to encode a spatial index with the size $D/2$, and a time index with the size $D/2$. with causal masks to implement the function, coined trajectory attention $\mathtt{TA}(\cdot;\cdot)$,
@@ 3.3 Feed-forward refinement
Although MMP architecture does not constrain the number of frames $W$, the finite memory of the system can pose a practical limit. When processing tens or hundreds of frames as the prediction horizon, a joint processing of whole frames can be impossible. In order to overcome the limitation, we introduce a feed-forward refinement technique to deal with prediction horizon beyond a chosen $W$. Specifically, when processing the tokens of an extra frame, e.g., $\tilde{\mathbf{E}}_{l}^{i|j}$, where $i\leq W$ and $j>W$, we exploit the pre-computed key and value tensors of $\tilde{\mathbf{E}}_{l}^{i|\{k<W\}}$, which we illustrate in Figure˜3. Since we train the model with the causal attention masking applied to the trajectory attention, these key and value tensors remain equivalent to the case where a larger input size were considered to include the extra frame.
@@ 4 Experiment
In this section, we present the experimental details and compare MMP to state-of-the-art baselines. In Sections˜4.1 and 4.2, we provide the training details, the data processing, and the evaluation protocols. Then, we experiment, in Section˜4.3, the feed-forward prediction of pointmaps, and the ablation study in Section˜4.4.
@@ 4.1 Training details
Despite our design to maintain the strong 3D prior of the pre-trained model (Wang et al., 2024), the synthetic scenes in the training dataset can cause a distribution shift in visual texture. Therefore, we test the trade-off between different training schedules for mixing the synthetic and the real frames, then choose the default setting that demonstrates a balanced performance (see Section˜4.4 for more details). Our default setting trains MMP for 30 epochs using the AdamW optimizer (Loshchilov and Hutter, 2019) with 20k clips of length $W=5$ per epoch, the mini-batch size 16, and the learning rate $1\times 10^{-4}$. We sample the clips from real scenes for the first 5 epochs, then employ synthetic scenes for the rest of the training steps.
@@ 4.2 Evaluation details
To evaluate the feed-forward predictions (Section˜4.3), we employ 3 different test datasets covering dynamic scenes: Point Odyssey (Zheng et al., 2023), Sintel (Butler et al., 2012), and iPhone dataset (Gao et al., 2022). Point Odyssey and Sintel are synthesized scenes generated using 3D rendering engines (Zheng et al., 2023; Butler et al., 2012), and iPhone dataset covers real scenes captured using a synchronized set of camera, lidar, and IMU sensors (Gao et al., 2022). For each scene, we consider overlapping slices of 12 frames as the evaluation samples. [^2] We also downsample iPhone dataset (Gao et al., 2022) to 3fps to promote a larger motion. We measure the regression accuracy of the pointmaps predicted by MMP and the baselines: DUSt3R (Zhang et al., 2025), Robust-CVD (Kopf et al., 2021), CasualSAM (Zhang et al., 2022), and state-of-the-art MonST3R (Zhang et al., 2025). Specifically, using a strided sampling, we experiment with $W=2$ (stride 6), $W=4$ (stride 3), and $W=6$ (stride 2) for inference. As for the metric, we employ the scale and shift invariant error provided by the open source repository of MonST3R (Zhang et al., 2025) and report the median error in the target pointclouds per setting: M@2, M@4, and M@6 in Table˜1.
@@ 4.3 Feed-forward pointmap prediction
In this section, we experiment the feed-forward pointmap prediction by MMP. In Table˜1, we quantitatively compare the quality of pointmap regression by MMP and the baselines: DUSt3R (Zhang et al., 2025), Robust-CVD (Kopf et al., 2021), CasualSAM (Zhang et al., 2022), and MonST3R (Zhang et al., 2025). Next, we provide the visualization of the pointmaps produced by MMP in Figure˜4, executed on DAVIS video frames (Perazzi et al., 2016).
To begin with, we find MMP can outperform the strongest feed-forward baseline, MonST3R (Zhang et al., 2025), e.g., 15.1% improvement M@6 1.772 (MonST3R (Zhang et al., 2025)) $\to$ 1.504 (MMP) on iPhone dataset (Gao et al., 2022) in Table˜1. While our method is trained on the same data distribution as the baseline, an enhanced performance is observed even under a pair-wise inference (i.e., M@2). This supports the significance of the trajectory encoder employed in our method, which facilitates learning useful representation for predicting accurate pointmaps. MMP can consistently improve the quality of dynamic pointmaps compared to the baselines in various scenarios covering synthetic and real video scenes. We also note that our method can demonstrate the results that are more robust over various strides, (e.g., comparing to MonST3R (Zhang et al., 2025) in Sintel (Butler et al., 2012): 1.374 $\to$ 1.298 M@2, 1.411 $\to$ 1.288 M@4, and 1.433 $\to$ 1.287 M@6), which we attribute to the dynamics modeling enabled by our method.
@@
From the qualitative study in Figure˜4, we find our method tends to demonstrate more accurate pointmaps over the frames, e.g., the background objects and the scene are consistently depicted, comparing the regions indicated by red boxes (MonST3R (Zhang et al., 2025)) and the blue boxes (MMP), which reveals the efficacy of our method in complex dynamic scenes.
@@ 4.4 Ablation study
We further study the effect of applying different schedules for training synthetic and real scenes by MMP in Table˜3, which compares 4 different training strategies: synthetic only (i.e., synthetic scenes for 30 epochs), joint training (i.e., mixed data for 30 epochs), synthetic then real (i.e., synthetic scenes for the first 25 epochs, then real scenes for the rest 5 epochs), and real then synthetic (the default setting). While there exist trade-offs in the performances over the datasets, we choose the real then synthetic schedule as our final design, which can demonstrate a balanced performance.
@@ 5.1 Extreme case
However, in practice, including the scenarios represented by our evaluation, there exists an extreme case where $\mathbf{I}^{i}$ and $\mathbf{I}^{k}$ are completely non-overlapping, so that the pair-wise model assigns an erroneous estimate of $\mathtt{Pr}(\mathbf{Y}^{i|k})$, which can induce significant failure modes of estimating the joint density. Even if the global optimization is employed, depending on the sampling strategy, there is a potential extreme case that the connectivity becomes independent. To prevent the case, a sophisticated hyperparameter engineering would be required. Since MMP can relax this constraint up to $W$ frames and beyond (with the feed-forward refinement), it can learn the pointmap distribution that is more close to the true nature of the dynamic scenes. For example, the intriguing tendency of MMP in Table˜1, being robust to the evaluation stride can be attributed to a more accurate estimation of the joint density over a set of frames.
@@ 5.2 Limitation
Despite the promising results demonstrated by MMP, the scarcity of 4D dyanamic scenes can hinder the generalization performance. To mitigate the distribution shifts, we employ the scheduled training to maintain the visual texture prior in the pre-trained model. However, since we still observe trade-offs in the performance, as shown in Table˜3, designing new training datasets, self-supervised learning with unlabeled data, or an objective functions robust to the distribution shift for 4D geometry estimation can be interesting future directions. It is also worth noting that we focus on the realistic scenarios where the observation is captured by a monocular video camera, rather than multiple synchronized cameras capturing one scene. Although it would be straightforward to apply MMP for the synchronized cameras, we believe that there is a room to exploit useful properties, such as epipolar geometry (Hartley and Zisserman, 2003) of the synchronized cameras, which is another interesting future direction.